OpenStudy (anonymous):

PLEASE HELP ME Simplify the expression. 2^3 * 2^5 A. 4^15 B. 4^8 C. 2^15 D. 2^8

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

\(\Large\sf a^b \times a^c = a^{b+c}\)

OpenStudy (anonymous):

B?

TheSmartOne (thesmartone):

Notice that 'a' You did it correctly by adding b+c But notice that the 'a' stays the same :)

OpenStudy (anonymous):

so 4^15

OpenStudy (anonymous):

or I had it right the first time?

TheSmartOne (thesmartone):

\(\Large\sf a^b \times a^c = a^{b+c}\) \(\sf a= 2\\ b= 3\\ c= 5\)

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

Its not B ._.

TheSmartOne (thesmartone):

Notice the A doesn't change

OpenStudy (anonymous):

no its A because 3 * 5

TheSmartOne (thesmartone):

But we are adding the exponents @ybnjgbnjgbn

TheSmartOne (thesmartone):

\(\Large\sf a^b \times a^c = a^{b+c}\) \(\sf a= 2\\ b= 3\\ c= 5\) Plug in the numbers and tell me what you get :)

OpenStudy (anonymous):

D because we add the same and we add 3 + 5

OpenStudy (anonymous):

its D @TheSmartOne

TheSmartOne (thesmartone):

Yes :D :D :D

TheSmartOne (thesmartone):

D is correct :)

OpenStudy (anonymous):

thanks!!